OpenStudy (anonymous):

Can someone help me?? Consider the following pair of equations: y = –2x + 8 y = x – 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (unklerhaukus):

substituting y –2x + 8 = x – 1 solve this for x then substitute x back into either of the original equations, to solve for y

OpenStudy (anonymous):

So is this correct then?? –2x + 8 = x – 1 +2x +2x 8 = 3x - 1 +1 +1 9 = 3x 3 = x -2(3) + 8 = 3 - 1 -6 + 8 = 3 - 1 2 = 2

OpenStudy (unklerhaukus):

you have found x =3, that's right! now, to find y : sub x=3 into one of the original equations (say the second one) y = (3) – 1 =